Brenot Corporation uses activity-based costing to determine product costs for external financial reports. Activity rates computed at the beginning of the year are used to apply manufacturing overhead costs to products. The company has provided the following data concerning its activity-based costing system.
The data used to develop activity rates were:

Activity Cost Pools

Estimated Overhead Cost

Expected Activity

Machine related (machine-hours)

$379,600

13,000

MHs

Batch setup (setups)

$1,144,800

27,000

setups

General factory (direct labor-hours)

$420,500

29,000

DLHs

The actual activity for the year was:

Activity Cost Pools

Total

Product X

Product Y

Machine related (machine-hours)

13,000

11,000

2,000

Batch setup (setups)

26,000

3,000

23,000

General factory (direct labor-hours)

30,000

24,000

6,000

The actual total manufacturing overhead cost incurred for the year was $1,942,300.

The activity rate computed at the beginning of the year for the batch setup activity cost pool is closest to:

Select one:

a. $42.40

b. $74.80

c. $49.80

d. $381.60

The total amount of overhead cost allocated to Product X during the year would be closest to:

Select one:

a. $1,145,000

b. $972,450

c. $796,400

d. $224,000

The credits to the Manufacturing Overhead control account during the year (prior to closing out the balance) would have totaled:

Select one:

a. $1,932,300

b. $1,929,650

c. $1,917,000

d. $1,942,300

The debits to the Manufacturing Overhead control account during the year (prior to closing out the balance) would have totaled:

Select one:

a. $1,932,200

b. $1,929,650

c. $1,942,300

d. $1,917,000
Answer #1

1) Activity rate of batch setup = 1144800/27000 = 42.40 per setup

So answer is a) $42.40

2) Total amount of overhead cost allocated to product X

Product x
Machine related 379600/13000*11000 = 321200
Batch setup 3000*42.40 = 127200
General factory 420500/29000*24000 = 348000
Total 796400

c) $796400

3) Credit to manufacturing overhead control account

Machine related 379600
Batch setup (42.40*26000) 1102400
General factory (14.50*30000) 435000
Total 1917000

So answer is c) $1917000

4) Debit to manufacturing overhead control account = Actual overhead = 1942300

So answer is c) $1942300